tan theta + cot theta= 2

Grade:11

3 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ \\tan\theta+{ 1 \over tan\theta}=2 \\=>tan\theta=1 \\=>\theta=45^0 \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Sahil suyal
5 years ago
Let us denote theta as β
So, 
Tanβ+cotβ=2,we need to prove it
As we know 
Cotβ=1/tanβ
So, 
Tanβ+1/tanβ=2
This would imply that
Tanβ=1
Now, 
Tanβ=Tan45
Implying 
β=45
This is what is you get as your answer
 
 
By Sahil suyal
Student
ankit singh
3 years ago
tan2theta+1-2tantheta=0
there for by shridharacharya formula gensolution is npi+-pi/2
in first quadrant theta=45